# Creativity and Machine Learning: A Survey

### v4

@@ methods
As far as the type of creativity is concerned, there can be ways to achieve a better exploration or even transformation of the space of solutions. For example, since CAN novelty loss is used during training, it learns to diverge from the distribution of real data. The same is true for RL-based methods with novelty and surprise rewards (especially if the training happens from scratch). Finally, a more explored or transformed space may be reached using RL-based methods driven by curiosity (Burda et al., 2019): an agent can learn to be creative and discover new patterns thanks to intrinsic rewards to measure novelty, interestingness and surprise. This can be done by training a predictive model of the growing data history and by using its learning progress as the reward. In this way, the agent is motivated to make things the predictor does not yet know. If an external qualitative reward is considered as well, the agent should in theory learn to do things that are new, but still valuable (Schmidhuber, 2010). The same idea can also be applied to different techniques like evolutionary strategies (Machado et al., 2007). Deep Learning Novelty Explorer (DeLeNoX) (Liapis et al., 2013a) uses a denoising autoencoder to learn low-dimensional representations of last generated artifacts. Then, a population of candidate artifacts (in terms of their representation) is evolved through a feasible-infeasible novelty search (Liapis et al., 2013b) in order to maximize the distances between them, i.e., to increase their novelty, while still considering qualitative constraints. Other evolutionary strategies might be considered as well to search the space of artifacts for novel (Lehman and Stanley, 2011) and surprising (Gravina et al., 2016) results. Instead of relying on manually crafted metrics, Quality Diversity through Human Feedback (QDHF) (Ding et al., 2023) uses human feedback for computing quality and distance in learned latent projection for computing diversity. Quality-Diversity through AI Feedback (QDAIF) (Bradley et al., 2023) makes the model more independent in searching and innovating by totally relying on its own feedback for both quality and diversity.
@@ 4. Creativity Measures
In this section, we present different methodologies to evaluate the creativity of artifacts generated by artificial agents. These can typically be extended to human-generated artifacts. For each of them, we explore the core concepts, the dimensions of creativity that are considered, the evaluation protocol, and, finally, we critically assess them.
The presence of several different proposals can be associated to the fact that it is not always straightforward to determine the “right” question to ask in an evaluation of a creative artifact (Gatt and Krahmer, 2018). For instance, the Generation, Evaluation, and Metrics (GEM) Benchmark for natural language generation (Gehrmann et al., 2021) does not contain any creativity evaluation measure. This is due to the inadequacy of the creative metrics proposed to date to correctly capture and measure all the necessary dimensions of creativity. The focus of our overview is on measures that are based on (or associated with) machine learning techniques. It is worth noting that some of them might be calculated without using machine learning, but we will refer to an implementation based on the latter. For an in-depth overview about creativity measures not strictly related to machines, we refer to (Said Metwaly et al., 2017).
Table 2 reports all the evaluation methods considered in this section, highlighting the dimensions they try to capture, their applicability and limitations. We will discuss these aspects in the remainder of this section.
Name What evaluates How evaluates Applicability
Limits
@@
What evalua-
tors state
creativity is
@@
    No chance to correctly com-
pute creativity for last works
@@
Regent-Dependent
creativity metric
Novelty,
value
Bayesian surprise and
synergy
General
    Requires to define and
extract features
@@ 4.1.1. Overview
The Lovelace Test (LT) (Bringsjord et al., 2001) was proposed in 2001 as a creativity-oriented alternative to the world famous Turing test (Turing, 1950). More formally, LT is defined as follows:
Definition 4.1 ().
@@
2) $A$’s outputting $o$ is not the result of a fluke hardware error, but rather the result of processes $A$ can repeat;
@@ 4.1.2. Dimensions of Creativity Considered
Since the evaluation depends on the tests performed by human evaluators, the dimensions of creativity considered by it might vary greatly. This allows for considering value, novelty and surprise, as well as domain-specific dimensions.
@@ 4.1.4. Critical Examination
This methodology represents an effective way to measure creativity, since it is ideally applicable to any field and it is quantitative. Although the latter is not based on machine learning, it may be used in principle for performing (some of) the tests. However, this methodology requires considerable human intervention in order to define all the tests.
@@ 4.2.1. Overview
Ritchie’s Criteria (Ritchie, 2007) are a set of criteria for evaluating the extent to which a program has been creative (or not) in generating artifacts. These criteria are based on three main factors: novelty, quality, and typicality. (Ritchie, 2007) contains a proposed series of criteria, but according to the authors they should only be intended as a “repertoire”.
@@ 4.3.1. Overview
In (Colton et al., 2011) the authors introduce FACE as a descriptive model of the creative act of artificial systems. A creative act is considered as a non-empty tuple of generative acts. FACE is designed to provide assessors with both quantitative and qualitative evaluations.
@@ 4.3.3. Protocol for Evaluation
More specifically, the FACE model considers a creative act as a non-empty tuple of generative acts of eight possible types: an expression of a concept, i.e., an instance of an (input, output) pair produced by the system; a method for generating expressions of a concept; a concept; a method for generating concepts; an aesthetic measure; a method for generating aesthetic measures; an item of framing information, i.e., an additional information or description regarding the generative act; a method for generating framing information. It is then possible to use it in a quantitative way (how many acts are produced); in a cumulative way (how many types of acts are considered); and in a qualitative way (by means of the aesthetic measure taken into consideration).
@@ 4.4.3. Protocol for Evaluation
SPECS is composed of three steps. The first one requires to provide a definition of creativity that the system should satisfy, using the suggested components, and potentially other domain-specific ones. The second requires to specify the standards to be used to evaluate such components. Finally, the third requires to test the system against the standards and report the results.
@@ 4.4.4. Critical Examination
This is an effective framework for working with computational creativity, but it cannot be considered as a practical evaluation method. Its effectiveness is strongly dependent on which components are considered and how they are evaluated for each specific task. In (Jordanous, 2014) the author discusses how SPECS satisfies Ritchie’s criteria and it is more comprehensive and expressive than the FACE model and the creative tripod (Colton, 2008) (according to which a creative system should exhibit skills, appreciation, and imagination) in a meta-evaluation test. They also use a human opinion survey based on five criteria: correctness, usefulness, faithfulness (as a model of creativity), usability (of the methodology), and generality. SPECS is evaluated considering music improvisation generators; it is also judged by the developers of the generative systems. In particular, the authors show that SPECS can help obtain additional insights on how a generative model works and how it can be improved.
@@ 4.5.4. Critical Examination
The Creativity Implication Network represents an effective way to deal with the creativity of sets of artworks. It considers both value and novelty, and it allows for using automated techniques in the computation of similarity. On the other hand, two potential drawbacks should be highlighted. The first one is related to artworks that occupy the position of “leaves” in the graph: if there are no subsequent works in the graph, their creativity would only be based on novelty, and not on value. The second one is more subtle, and it is about time-positioning. As demonstrated by (Elgammal and Saleh, 2015), moving back an artwork has the effect of increasing its creativity; however, this appears as conceptually wrong. As discussed in (Boden, 2003), the time location of an artwork is fundamental in the quantification of its creativity. It may happen that, due to the surprise component of creativity, an artwork that appears too early might not be considered as surprising, because observers are not able to truly understand it; on the contrary, if it appears too late, it might be considered as obvious and not surprising at all. In conclusion, even if this approach is able to correctly capture value and novelty, it cannot capture the concept of surprise.
@@ 4.6.1. Overview
Generate-and-test setting (Toivonen and Gross, 2015) is a family of methods based on the separation of the generative process in two phases: generation and evaluation. Let us assume that gen() denotes the generative function of the system, and a an artifact generated by the function. Then, a is evaluated by another function eval() that assesses the degree of creativity of a. If it passes the evaluation, it is outputted by the system. For example, the authors of (Varshney et al., 2019) use this approach to develop a computational creativity system for generating culinary recipes and menus called Chef Watson. (Colton et al., 2015) describes an augmentation of Painting Fool (Colton, 2012) with Digital ARtist Communicating Intention (DARCI) (Norton et al., 2010) in a generate-and-test setting (i.e., by using The Painting Fool for generation, and DARCI for evaluation). Pseudo-Intelligent Evolutionary Real-time Recipe Engine (PIERRE) (Morris et al., 2012) is based as well on two models, one for generating recipes by means of a genetic algorithm, and one for evaluating them.
@@ 4.6.3. Protocol for Evaluation
The evaluation protocol strictly depends on the specific implementation of the eval() function. For instance, Chef Watson (Varshney et al., 2019) uses two measures: flavorfulness for quality, and Bayesian surprise for novelty. Flavorfulness is computed by means of a regression model, built on olfactory pleasantness considering its constituent ingredients. Bayesian surprise (Baldi and Itti, 2010) is a measure of surprise in terms of the impact of a piece of data that changes a prior distribution into a posterior distribution, calculated applying Bayes’ theorem. The surprise is then the distance between the posterior and prior distributions. It is worth noting that it has been demonstrated that there exists a mathematical limit in the maximization of quality and novelty when novelty is expressed in terms of Bayesian Surprise (Varshney, 2019).
On the other hand, The Painting Fool (Colton, 2012) uses DARCI (Norton et al., 2010) in place of the evaluation function. DARCI is able to make associations between image features and descriptions of the images learned using a series of NNs as the basis for the appreciation. It has therefore been used as a sort of artificial art critic to complement The Painting Fool, allowing it to assess the validity of its own creations.
Finally, PIERRE (Morris et al., 2012) evaluates the generated recipes again using novelty and quality. Novelty is computed by counting new combinations of ingredients used. Quality is based on two NNs that perform a regression of user ratings based on the amount of different ingredients, working at two levels of abstraction.
@@ 4.7.1. Overview
Even if not strictly related to creativity, (Burns, 2006) proposes an approach to measure aesthetic experience called Atoms of EVE’, which is based on a probabilistic model of the world to derive expectations and explanation. The authors state that aesthetic arises in two ways: by forming E (expectation) while avoiding V (violation); and by forming E’ (explanation) while resolving V.
@@ 4.7.4. Critical Examination
As observed before, while such a computation for surprise is common, this is not true for value. The focus on explanation provides an interesting way to mathematically define value. However, value is not only about finding meaning, but also about utility, performance and attractiveness (Maher, 2010); this is not possible through this measure. Finally, novelty is not considered at all.
@@ 4.8.1. Overview
The authors of (Maher and Fisher, 2012) propose a model to evaluate creativity in the specific domain of design. They consider creativity as a relative measure in a conceptual space of potential and existing designs. Designs are represented by attribute-value pairs; and novelty, value and surprise capture different aspects of creativity in their space.
@@ 4.8.3. Protocol for Evaluation
The model is based on the analysis of the conceptual space of potential and existing designs defined by all the potential attribute-value pairs. Novelty is evaluated with respect to a description space, i.e., by considering each product as the set of its descriptive attributes. Value is considered with respect to a performance space, i.e., by considering attributes that have utility preferences associated to them. Finally, surprise is based on finding violations of patterns that are possible to anticipate in the space of both current and possible designs. The K-means clustering algorithm is used to organize known designs by means of their attributes. Then, novelty, value and surprise measures of a new design are obtained by looking at the distance to the nearest cluster centroid.
@@ 4.8.4. Critical Examination
Even if it explicitly targets the design domain, this approach is able to combine the three dimensions of creativity by Boden. Nonetheless, it is limited by the fact that artifacts have to be described through an attribute-value pair representation. In particular, a large number of features might be needed. Otherwise, we might lose aspects of the artifacts that are fundamental to correctly quantify creativity. Since it is not possible to know the fundamental features in advance, the method requires one to enumerate as many features as possible. However, the risk is to define an excessive number of non-informative attributes, making the computation of the metrics too computationally expensive. In fact, the data points become increasingly “sparse” as dimensionality increases; many techniques (especially clustering) are based on distance, and therefore they may suffer from the curse of dimensionality (Steinbach et al., 2004). Finally, as for classic machine learning techniques, there is the need of manually defining and extracting the chosen features from unstructured data, which is a time-consuming and potentially prone-to-error activity. A possible way to overcome the problems related to feature extraction and the curse of dimensionality might be to adopt deep learning techniques, given their effectiveness with unstructured data.
4.9. Regent-Dependent Creativity Metric
4.9.1. Overview
The Regent-Dependent Creativity metric (França et al., 2016) is based on novelty and value. The name comes from the Regent-Dependent Model, a proposed method to describe artifacts as collections of features, i.e., through sets of pairs with a regent (e.g., an action or an attribute) and a dependent (e.g., the specific target or value of the regent), labeled with a value to express the intensity of that pair in that context.
4.9.2. Dimensions of Creativity Considered
As explained above, the Regent-Dependent creativity metric considers novelty and value as dimensions of creativity. In particular, value is dependent on the associations and rules that bond artifacts in a specific context, which defines what is worth of interest and what is not.
4.9.3. Protocol for Evaluation
Novelty and value are computed through Bayesian surprise (see Section 4.6.3) and synergy, respectively. In particular, synergy is a metric that captures the cooperative nature of constituent elements of an artifact. It is based on modeling the artifacts as graphs. The vertices are the dependency pairs of the artifact. The edges between pairs exist only if they belong to the same set of synergistic variables, i.e., if they exhibit a better effect together than separately.
4.9.4. Critical Examination
In contrast to the approach described previously, the Regent-Dependent metric appears to be of more general applicability. However, only novelty and value are considered, and surprise is missing. In addition, just as detailed in Section 4.8.4, this method is limited by requiring the definition and extraction of all the features to compute creativity.
@@ 4.10.1. Overview
The authors of (Grace and Maher, 2014) suggest that a framework of unexpectedness (i.e., violation of an observer’s expectations) can deal with novelty, surprise and domain transformation (also called transformativity). Although they do not claim it can be a measure of creativity on its own, and that value should be added as well, they suggest it can become a vital component in computational creativity evaluation.

### v6

@@ methods
As far as the type of creativity is concerned, there can be ways to achieve a better exploration or even transformation of the space of solutions. For example, since CAN novelty loss is used during training, it learns to diverge from the distribution of real data. The same is true for RL-based methods with novelty and surprise rewards (especially if the training happens from scratch). Finally, a more explored or transformed space may be reached using RL-based methods driven by curiosity (Burda et al., 2019): an agent can learn to be creative and discover new patterns thanks to intrinsic rewards to measure novelty, interestingness, and surprise. This can be done by training a predictive model of the growing data history and by using its learning progress as the reward. In this way, the agent is motivated to make things the predictor does not yet know. If an external qualitative reward is considered as well, the agent should in theory learn to do things that are new, but still valuable (Schmidhuber, 2010). The same idea can also be applied to different techniques like evolutionary strategies (Machado et al., 2007). Deep Learning Novelty Explorer (DeLeNoX) (Liapis et al., 2013a) uses a denoising autoencoder to learn low-dimensional representations of the last generated artifacts. Then, a population of candidate artifacts (in terms of their representation) is evolved through a feasible-infeasible novelty search (Liapis et al., 2013b) in order to maximize the distances between them, i.e., to increase their novelty, while still considering qualitative constraints. Other evolutionary strategies might be considered as well to search the space of artifacts for novel (Lehman and Stanley, 2011) and surprising (Gravina et al., 2016) results. Instead of relying on manually crafted metrics, Quality Diversity through Human Feedback (QDHF) (Ding et al., 2023) uses human feedback for computing quality and distance in learned latent projection for computing diversity. Quality-Diversity through AI Feedback (QDAIF) (Bradley et al., 2024) makes the model more independent in searching and innovating by totally relying on its own feedback for both quality and diversity.
@@ 4. Creativity Measures
In this section, we present different methodologies to evaluate the creativity of artifacts generated by artificial agents. These can typically be extended to human-generated artifacts. For each of them, we explore the core concepts, the dimensions of creativity that are considered, the evaluation protocol, and, finally, we critically assess them. The presence of several different proposals can be associated with the fact that it is not always straightforward to determine the “right” question to ask in an evaluation of a creative artifact (Gatt and Krahmer, 2018). For instance, the Generation, Evaluation, and Metrics (GEM) Benchmark for natural language generation (Gehrmann et al., 2021) does not contain any creativity evaluation measure. This is due to the inadequacy of the creative metrics proposed to date to correctly capture and measure all the necessary dimensions of creativity. The focus of our overview is on measures that are based on (or associated with) machine learning techniques. It is worth noting that some of them might be calculated without using machine learning, but we will refer to an implementation based on the latter. For an in-depth overview of creativity measures not strictly related to machines, we refer to (Said Metwaly et al., 2017). Table 2 reports all the evaluation methods considered in this section, highlighting the dimensions they try to capture, their applicability, and their limitations. We will discuss these aspects in the remainder of this section.
Name What evaluates How evaluates Applicability Limits
@@
Evaluators’
creativity definition
@@
  Not possible to accurately
measure the creativity of
the most recent works
@@ 4.1.1. Overview
The Lovelace Test (LT) (Bringsjord et al., 2001) was proposed in 2001 as a creativity-oriented alternative to the world-famous Turing test (Turing, 1950). More formally, LT is defined as follows:
Definition 4.1.
@@
2) $A$’s outputting $o$ is not the result of a fluke hardware error but of processes $A$ can repeat;
@@ 4.1.2. Dimensions of Creativity Considered
Since the evaluation depends on the tests performed by human evaluators, the dimensions of creativity considered by it might vary greatly. This allows for considering value, novelty, and surprise, as well as domain-specific dimensions.
@@ 4.1.4. Critical Examination
This methodology represents an effective way to measure creativity since it is ideally applicable to any field and it is quantitative. Although the latter is not based on machine learning, it may be used in principle for performing (some of) the tests. However, this methodology requires considerable human intervention in order to define all the tests.
@@ 4.2.1. Overview
Ritchie’s Criteria (Ritchie, 2007) is a set of criteria for evaluating the extent to which a program has been creative (or not) in generating artifacts. These criteria are based on three main factors: novelty, quality, and typicality. (Ritchie, 2007) contains a proposed series of criteria, but according to the authors they should only be intended as a “repertoire”.
@@ 4.3.1. Overview
In (Colton et al., 2011) the authors introduce FACE as a descriptive model of the creative act of artificial systems. A creative act is considered a non-empty tuple of generative acts. FACE is designed to provide assessors with both quantitative and qualitative evaluations.
@@ 4.3.3. Protocol for Evaluation
More specifically, the FACE model considers a creative act as a non-empty tuple of generative acts of eight possible types: an expression of a concept, i.e., an instance of an (input, output) pair produced by the system; a method for generating expressions of a concept; a concept; a method for generating concepts; an aesthetic measure; a method for generating aesthetic measures; an item of framing information, i.e., a piece of additional information or description regarding the generative act; a method for generating framing information. It is then possible to use it in a quantitative way (how many acts are produced); in a cumulative way (how many types of acts are considered); and in a qualitative way (by means of the aesthetic measure taken into consideration).
@@ 4.4.3. Protocol for Evaluation
SPECS is composed of three steps. The first one requires providing a definition of creativity that the system should satisfy, using the suggested components, and potentially other domain-specific ones. The second requires to specify the standards for evaluating such components. The third requires to test the system against the standards and report the results.
@@ 4.4.4. Critical Examination
This is an effective framework for working with computational creativity, but it cannot be considered as a practical evaluation method. Its effectiveness is strongly dependent on which components are considered and how they are evaluated for each specific task. In (Jordanous, 2014) the author discusses how SPECS satisfies Ritchie’s criteria and it is more comprehensive and expressive than the FACE model and the creative tripod (Colton, 2008) (according to which a creative system should exhibit skills, appreciation, and imagination) in a meta-evaluation test. They also use a human opinion survey based on five criteria: correctness, usefulness, faithfulness (as a model of creativity), usability (of the methodology), and generality. SPECS is evaluated considering music improvisation generators; it is also judged by the developers of the generative systems. In particular, the authors show that SPECS can help obtain additional insights into how a generative model works and how it can be improved.
@@ 4.5.4. Critical Examination
The Creativity Implication Network represents an effective way to deal with the creativity of sets of artworks. It considers both value and novelty, and it allows for using automated techniques in the computation of similarity. On the other hand, two potential drawbacks should be highlighted. The first one is related to artworks that occupy the position of “leaves” in the graph: if there are no subsequent works in the graph, their creativity would only be based on novelty, and not on value. The second one is more subtle, and it is about time-positioning. As demonstrated by (Elgammal and Saleh, 2015), moving back an artwork has the effect of increasing its creativity; however, this appears conceptually wrong. As discussed in (Boden, 2003), the time location of an artwork is fundamental in the quantification of its creativity. It may happen that, due to the surprise component of creativity, an artwork that appears too early might not be considered as surprising because observers are not able to truly understand it; on the contrary, if it appears too late, it might be considered as obvious and not surprising at all. In conclusion, even if this approach is able to correctly capture value and novelty, it cannot capture the concept of surprise.
@@ 4.6.1. Overview
Generate-and-test setting (Toivonen and Gross, 2015) is a family of methods based on the separation of the generative process into two phases: generation and evaluation. First, the system generates a candidate artifact. Then, it evaluates its degree of creativity and outputs the artifact if the evaluation is passed. For example, the authors of (Varshney et al., 2019) use this approach to develop a computational creativity system for generating culinary recipes and menus called Chef Watson. (Colton et al., 2015) describes an augmentation of Painting Fool (Colton, 2012) with Digital ARtist Communicating Intention (DARCI) (Norton et al., 2010) in a generate-and-test setting (i.e., by using The Painting Fool for generation, and DARCI for evaluation). Pseudo-Intelligent Evolutionary Real-time Recipe Engine (PIERRE) (Morris et al., 2012) is also based on two models, one for generating recipes with a genetic algorithm, and one for evaluating them.
@@ 4.6.3. Protocol for Evaluation
The evaluation protocol strictly depends on the specific implementation of the eval() function. For instance, Chef Watson (Varshney et al., 2019) uses two measures: flavorfulness for quality, and Bayesian surprise for novelty. Flavorfulness is computed by means of a regression model, built on olfactory pleasantness considering its constituent ingredients. Bayesian surprise (Baldi and Itti, 2010) is a measure of surprise in terms of the impact of a piece of data that changes a prior distribution into a posterior distribution, calculated applying Bayes’ theorem. The surprise is then the distance between the posterior and prior distributions. It is worth noting that it has been demonstrated that there exists a mathematical limit in the maximization of quality and novelty when novelty is expressed in terms of Bayesian surprise (Varshney, 2019). On the other hand, The Painting Fool (Colton, 2012) uses DARCI (Norton et al., 2010) in place of the evaluation function. DARCI is able to make associations between image features and descriptions of the images learned using a series of NNs as the basis for the appreciation. It has therefore been used as a sort of artificial art critic to complement The Painting Fool, allowing it to assess the validity of its own creations. Finally, PIERRE (Morris et al., 2012) evaluates the generated recipes again using novelty and quality. Novelty is computed by counting new combinations of ingredients used. Quality is based on two NNs that perform a regression of user ratings based on the amount of different ingredients, working at two levels of abstraction.
@@ 4.7.1. Overview
(Burns, 2006) proposes an approach to measure aesthetic experience called Atoms of EVE’, which is based on a probabilistic model of the world to derive expectations and explanations. The authors state that aesthetic arises in two ways: by forming E (expectation) while avoiding V (violation); and by forming E’ (explanation) while resolving V.
@@ 4.7.4. Critical Examination
As observed before, while such a computation for surprise is common, this is not true for value. The focus on explanation provides an interesting way to mathematically define value. However, value is not only about finding meaning but also about utility, performance, and attractiveness (Maher, 2010); this is not possible through this measure. Finally, novelty is not considered.
@@ 4.8.1. Overview
The authors of (Maher and Fisher, 2012) propose a model to evaluate creativity in the specific domain of design. They consider creativity as a relative measure in a conceptual space of potential and existing designs. Designs are represented by attribute-value pairs; and novelty, value, and surprise capture different aspects of creativity in their space. A similar approach is at the basis of the regent-dependent model (França et al., 2016), according to which artifacts are described through sets of pairs with a regent (e.g., an action or an attribute) and a dependent (e.g., the specific target or value of the regent).
@@ 4.8.3. Protocol for Evaluation
The model is based on the analysis of the conceptual space of potential and existing designs defined by all the potential attribute-value pairs. Novelty is evaluated with respect to a description space, i.e., by considering each product as the set of its descriptive attributes. Value is considered with respect to a performance space, i.e., by considering attributes that have utility preferences associated to them. Finally, surprise is based on finding violations of patterns that are possible to anticipate in the space of both current and possible designs. The K-means clustering algorithm is used to organize known designs by means of their attributes. Then, novelty, value, and surprise measures of a new design are obtained by looking at the distance to the nearest cluster centroid.
@@ 4.8.4. Critical Examination
Even if it explicitly targets the design domain, this approach is able to combine the three dimensions of creativity by Boden. Nonetheless, it is limited by the fact that artifacts have to be described through an attribute-value pair representation. In particular, a large number of features might be needed. Otherwise, we might lose aspects of the artifacts that are fundamental to correctly quantify creativity. Since it is not possible to know the fundamental features in advance, the method requires one to enumerate as many features as possible. However, the risk is to define an excessive number of non-informative attributes, making the computation of the metrics too computationally expensive. In fact, the data points become increasingly “sparse” as dimensionality increases; many techniques (especially clustering) are based on distance, and therefore they may suffer from the curse of dimensionality (Steinbach et al., 2004). Finally, as for classic machine learning techniques, there is the need to manually define and extract the chosen features from unstructured data, which is a time-consuming and potentially prone-to-error activity. A possible way to overcome the problems related to feature extraction and the curse of dimensionality might be to adopt deep learning techniques, given their effectiveness with unstructured data.
@@ 4.9.1. Overview
The authors of (Grace and Maher, 2014) suggest that a framework of unexpectedness (i.e., violation of an observer’s expectations) can deal with novelty, surprise, and domain transformation (also called transformativity). Although they do not claim it can be a measure of creativity on its own, and that value should be added as well, they suggest it can become a vital component in computational creativity evaluation.